# On the Power Saving in High-Speed Ethernet-based Networks for Supercomputers and Data Centers

### v2

@@ Abstract
The increase in computation and storage has led to a significant growth in the scale of systems powering applications and services, raising concerns about sustainability and operational costs. In this paper, we explore power-saving techniques in high-performance computing (HPC) and datacenter networks, and their relation with performance degradation. From this premise, we propose leveraging Energy Efficient Ethernet (EEE), with the flexibility to extend to conventional Ethernet or upcoming Ethernet-derived interconnect versions of BXI and Omnipath.
We analyze the PerfBound proposal, identifying possible improvements and modeling it into a simulation framework. Through different experiments, we examine its impact on performance and determine the most appropriate interconnect. We also study traffic patterns generated by selected HPC and machine learning applications to evaluate the behavior of power-saving techniques.
From these experiments, we provide an analysis of how applications affect system and network energy consumption. Based on this, we disclose the weakness of dynamic power-down mechanisms and propose an approach that improves energy reduction with minimal or no performance penalty. To our knowledge, this is the first power management proposal tailored to future Ethernet-based HPC architectures, with promising results.
@@ 1 Introduction
Energy consumption in supercomputers and data centers is the sum of power drawn, among other elements, by computing and storage nodes, the interconnection network, and the cooling system. Considering the most powerful supercomputers, the Exascale barrier was first overrun by the Frontier supercomputer, which remained the first in the TOP500 ranking [top500] since June 2022, until El Capitan surpassed it in November 2024. Table 1 contains five of the most powerful supercomputers (in terms of computing power) with energy efficiency information supplied as per the last ranking. Energy efficiency is the ratio between the maximum performance and the average power consumption; the higher it is, the better. As we can see, the highest energy efficiency is achieved by the JUPITER booster machine, which is #4 in the ranking as of June 2025.
@@
The operational costs for supercomputers are concerning because of power availability and sustainability reasons. Green500 [green500] ranks TOP500 systems according to their power efficiency rather than the raw performance. If we examine the Green500 ranking for that date, we can see that the aforementioned systems are, in terms of power efficiency, in ranks #13, #42, #68, and #12, respectively. Of course, designing a system with high regard for power efficiency leads to lower operational costs, namely in electricity.
Processors and accelerators used in the most powerful supercomputers have special mechanisms for reducing power consumption when thermal or power constraints are present. Intel processors, for example, introduced the C-states in the Sandy Bridge architecture [intelsandybridge] for advanced power management techniques. CPUs have several states defined, providing lower power consumption at the cost of performance and latency. Operating systems can indicate to the processor to enter a lower power state. However, changing states involves microarchitectural operations and voltage transitions, increasing latencies. During state transition times, CPUs execute no code, affecting performance. Still, even the most energy-saving states consume power. These features aim to keep CPU energy consumption proportional to the usage.
@@
In this context, as computing nodes become more energy proportional, the interconnection network can become a significant fraction of computing system power [energyproportional]. Therefore, it is also necessary to design high-performance interconnection networks whose energy consumption is proportional to the traffic they carry and the rest of the system elements (e.g., processors, storage, etc).
@@
Mainly, the opportunities for power savings come from periods of inactivity that would allow us to set ports to power-saving states. This would make power consumption more proportional to the network usage. If we look at the network usage for several applications, we can see that network activity fluctuates during execution. Periods of low network usage are good opportunities for applying power-saving techniques. Figure 1 shows the inactivity periods for a specific port on several applications when executed on our target simulated environment. The data is then used to form a histogram of 200 fixed-sized bins. Each bin $i$ stores the number of registered inactivity periods that last between $width_{bin}\times i$ and $width_{bin}\times(i+1)$ seconds. In these plots, we show values up to the 99th percentile for visibility reasons. On the $x$ axis, we show the length of an inactivity period on that port. The left $y_{1}$ axis in black shows how many periods of inactivity are within the specified bin, and on the right, $y_{2}$ in red shows the cumulative distribution for every bin. In other words, the percentage of the histogram “area” below the bin. Regardless of the distribution on inactivity periods, any power-saving mechanism should take advantage of the longer periods to provide as much energy savings as possible. Moreover, we expect the performance penalty of the mechanism to be as low as possible or within a reasonable margin.
@@
This paper explores the effects of network power management strategies on the performance of interconnection networks. First, we describe the background concepts in network energy reduction (see Section 2). Next, we describe the network power for large Ethernet-based interconnection networks used in supercomputers and data centers, and how it can be applied to BXIv3-based networks (see Section 3). We also show that more efficient interconnection networks are required to achieve energy proportionality. Afterwards, we describe the power model for BXIv3 that we have designed and the proposed energy-saving strategies. This model has been implemented in a powerful simulation framework, allowing large-scale network simulations. Furthermore, we have performed a set of simulation experiments (see Section 4) that show that, with the proposed power model strategies, the network power can be lowered without gravely affecting the overhead on execution time or latency. In particular, we managed to save more than 5 % of energy on the network by focusing on links only, while still outperforming other previous proposals in terms of performance overhead. We manage to improve the PerfBound [perfbound] technique, even to the point of reverting power increases and resulting in actual savings. In our opinion, this is a promising start, given that our proposal and assumptions could be portable to future interconnection network technologies, such BXIv3 and the upcoming (Formerly Intel’s) Omnipath [^1] https://www.theregister.com/2025/06/09/omnipath_is_back/ [omnipath].
Mainly, the opportunities for power savings come from periods of inactivity that would allow us to set ports to power-saving states. This would make power consumption more proportional to the network usage. If we look at the network usage for several applications, we can see that network activity fluctuates during execution. Periods of low network usage are good opportunities for applying power-saving techniques. Figure 1 shows the inactivity periods for a specific port on several applications when executed on our target simulated environment. The data is then used to form a histogram of 200 fixed-sized bins. In these plots, we show values up to the 99th percentile for visibility reasons. On the $x$ axis, we show the length of an inactivity period on that port. The left $y_{1}$ axis in black shows how many periods of inactivity are within the specified bin, and on the right, $y_{2}$ in red shows the cumulative distribution for every bin. In other words, the percentage of histogram ”area” below the bin.
@@
In summary the main contributions of this paper are the following:
@@
    Study and implementation discussion of PerfBound [perfbound] in our network simulator.
@@ 2.1 Energy consumption in high-speed interconnection networks
Network performance is affected by several network design issues, such as the topology, routing, flow control, and operating frequency, among others. Thus, designing an efficient and cost-effective network is crucial to guarantee the performance of the entire supercomputer and datacenter, as well as a balance between this performance and the investment and operational costs. Regarding HPC network technologies, a quick look into the TOP500 list shows us that Gigabit Ethernet, in various revisions, is present in many systems (32.8 %), only second to Infiniband (54.2 %). Such is the momentum of Ethernet in these environments that technologies like iWarp [iwarp] and RoCE (RDMA over Converged Ethernet) [roceiwarp] are working towards providing Ethernet with RDMA, present in interconnects like Infiniband.
@@
Despite the notable capital investment for the equipment (CAPEX), operation costs (OPEX) and energy constraints are also relevant. The most notable examples of OPEX in HPC are power supply, cooling, and maintenance. Saving energy to reduce OPEX is being tackled by reducing power consumption (at the potential cost of performance), and cooling requirements are proportional to power consumption and thus, heat dissipation. The sheer scale of HPC environments implies huge CAPEX and OPEX. Electricity costs can be a deciding factor for HPC environments. Therefore, it is no surprise that power consumption reduction is sought, especially when examining the sparse network access in scientific applications. In fact, InfiniBand, the most pervasive interconnect in HPC, has power-saving mechanisms [dickov_software-managed_2014]. The ubiquity of Ethernet in other commercial areas has also caused several proposals to appear for Ethernet as well [gunaratne_reducing_2008, hays_activeidle_2008]. The IEEE Task Force assembled to provide the technology with a common standard, resulting in the standard IEEE 802.3az, also known as Energy Efficient Ethernet (EEE).
@@
Several design and operation strategies have been applied to reduce energy consumption in interconnection networks. Specifically, network energy consumption can be reduced by designing networks with high-radix switches. By increasing the switch radix, fewer routers are required. As switches aggregate more components, this also reduces the number of hops [kim_microarchitecture_2005]. Consequently, the number of links in the network is reduced when the hop count is lower. Routing algorithms can also be relevant to energy consumption if additional links provide several paths between two nodes. Of course, these links will consume energy when unused.
On the other hand, operation cost (OPEX) in HPC environments increases as more network elements and nodes are added. Like with processors, network energy consumption could be reduced when the network is not being used. Periods of inactivity are determined by the communication pattern of applications. Scientific simulations attempt to minimize inter-node communication, with sparse bursts of network activity. On the other hand, in data centers, we can find deep learning applications, distributed storage, cloud computing, and data-intensive applications (e.g., search engines). The network load for this second set of applications depends on user requests, which need to be served as soon as possible to reduce perceived latency by end users [andujar_extending_2023].
Ethernet network links are traditionally always on, available to be used up to maximum speed. While on, ports frequently communicate to maintain alignment and synchronization. However, if links are switched to a power-saving state, energy could be saved over long network inactivity periods. Various techniques or strategies have been proposed to address this issue, some of which are described in the following sections.
@@ 2.2 Initial attempts for reducing link power consumption
The power consumption of any device can be divided into static and dynamic power. The later depends on clock frequency, supply voltage and transistor technology. Power consumption is roughly proportional to the cube of the frequency, which is promising for power saving. Dynamic Voltage and Frequency Scaling (DVFS) adjust voltage and frequency based on network utilization to save energy during idle periods and maintain performance under load. Implementations like Adaptive Supply Serial Links (ASSL) adjust supply voltage dynamically, though hardware complexity and additional circuitry such as voltage regulators are required [kim_adaptive_2002, shang_dynamic_2003]. This conflicts with the objectives of interconnection networks being simple.
Since network technologies have gone through several revisions and can usually work at previous-generation data rates, this was thought to be a good strategy for energy saving. Adaptive Link Rate (ALR) aims to accomplish this by lowering link rates during periods of low utilization without altering voltages or frequencies directly. Instead, links can negotiate lower speeds. Although it can significantly save power, switching link rates introduces latency due to MAC handshakes, PHY configuration, and resynchronization [gunaratne_reducing_2008], making implementation complex and limiting practical adoption. This is detrimental in HPC environments because it can turn the network into a bottleneck when it is more needed.
2.3 Low Power Idle (LPI)
As ALR proved inefficient, Low Power Idle (LPI) [hays_activeidle_2008] was proposed. LPI turns components off during inactivity periods instead of modifying the link frequency, which was more suitable. The corresponding states for links are Wake for ports active for transmission, and Sleep for inactive ports in a power-saving state. The underlying idea is that higher data rates provide the most energy-efficient transmission in terms of Joules per bit. Contrary to ALR, the switch between Wake and Sleep states occurs in microseconds instead of nanoseconds. Figure 2 shows these states and how they relate to packet transmissions. Note that ${t}_{w}$ corresponds to the time to switch the port from Sleep to Wake, and ${t}_{s}$ for the opposite transition. Specific values for these transition times on different Ethernet revisions are displayed in Table 2 [specifications_ieee_2009].
@@ 2.3 Low Power Idle (LPI)
This figure shows that the link state starts at Sleep, and at a certain time, it starts transitioning to its Wake state. After ${t}_{w}$, the link is active and able to send several packets. Once the link has finished the transmission, it returns to Sleep. Note that every packet burst happening while the link is in Sleep state spends ${t}_{w}$ waiting for the link to transition, regardless of the burst size. Then, longer messages are more efficient to transmit because fewer state transitions occur. On the contrary, short messages may have to wait for the link to transition for longer than the packet requires for transmission. This is detrimental to latency when the network is sparsely used, as is the case in most HPC environments. Moreover, during transitions, link power consumption is the same as when in the Wake state.
The IEEE 802.3az draft [specifications_ieee_2009] specifies values for ${t}_{w}$ and ${t}_{s}$ for several Ethernet specifications, as shown in Table 2. For a comparison, the time to send a 1500-byte frame has also been considered. As we can see, higher speeds reduce frame times, but increase Sleep and Wake transition times considerably. Therefore, LPI is suitable for large bursts followed by large periods of inactivity.
Protocol ${t}_{w}$ ${t}_{w}$ Frame time (1500B)
100Base-Tx (100 Mb/s) 30 µs 100 µ 120 µs
@@
10GBase-T (10 Gb/s) 448 µs 288 µs 1.2 µs
Table 2: Link transition times for different Ethernet versions (extracted from [specifications_ieee_2009]).
2.4 Energy Efficient Ethernet (EEE)
The IEEE 802.3az standard, also known as Energy Efficient Ethernet (EEE) [christensen_ieee_2010, specifications_ieee_2009], makes use of the Low Power Idle (LPI) mechanism to reduce energy consumption during idle periods on Ethernet links. This mechanism defines two states: Active and LPI, which involve transitioning between normal operation and a low-power mode. The EEE standard specifies only the basic behavior of LPI, leaving implementation details such as transition policies to the vendors.
When EEE development started, only Sleep and Wake (sometimes referred to as Idle and Active, respectively) were envisioned to provide power savings. Later enhancements to the LPI mechanism were proposed in subsequent amendments to the IEEE 802.3 standard. Specifically, the IEEE 802.3bj amendment, which targets 40 Gb/s and higher-speed Ethernet links, introduces refinements such as the Fast Wake state [noauthor_options_2012]. Also, the former Sleep state is renamed as Deep Sleep. The Fast Wake state addresses the latency drawbacks of the original LPI proposal by maintaining partial activity in PHY components, while Deep Sleep shuts more components down. This reduces wake transition times significantly, from 4.6 µs in Deep Sleep to around 250-500 ns in Fast Wake. However, this comes at the cost of reduced power savings, typically around 20–40% for Fast Wake compared to up to 90% in Deep Sleep [saravanan_powerperformance_2013]. Anyway, the criterion for designing transitions at the hardware is still left to the vendors.
Figure 3 illustrates how employing different low-power states affects link transition times and, consequently, the overall performance of data transmission in energy-aware systems.
@@ 2.4 Energy Efficient Ethernet (EEE)
The different link power levels are a basis for implementing LPI, included in the EEE standard [christensen_ieee_2010]. This has led to several proposals tailored for HPC so that end-to-end latency is not greatly increased by the overhead of turning links on [reviriego_performance_2009, ANDUJAR2023126]. The idea for HPC environments is to employ Fast Wake so that latency is not greatly affected by ports powering up. However, transitioning up as soon as packet transmission ends causes ports to spend more time transitioning than transmitting packets. Table 3 shows a summary of the available states in EEE, with their link power and transition time.
State Link power Transition time to Wake
Wake 100 % -
Fast Wake 60-80 % 250 - 500 ns
Deep Sleep 10 % 4.48 µs
Table 3: Summary of EEE link power states (based on [noauthor_options_2012]).

### v3

@@ Abstract
The increase in computation and storage has led to a significant growth in the scale of systems powering applications and services, raising concerns about sustainability and operational costs. In this paper, we explore power-saving techniques in high-performance computing (HPC) and datacenter networks, and their relation with performance degradation. From this premise, we propose leveraging Energy Efficient Ethernet (EEE) protocol, with the flexibility to extend to conventional Ethernet or upcoming Ethernet-derived interconnect versions of BXI and Omnipath.
We analyze the PerfBound power-saving mechanism, identifying possible improvements and modeling it into a simulation framework. Through different experiments, we examine its impact on performance and determine the most appropriate interconnect. We also study traffic patterns generated by selected HPC and machine learning applications to evaluate the behavior of power-saving techniques.
From these experiments, we provide an analysis of how applications affect system and network energy consumption. Based on this, we disclose the weakness of dynamic power-down mechanisms and propose an approach that improves energy reduction with minimal or no performance penalty. To the best of our knowledge, this work presents the first thorough analysis of PerfBound and an enhancement to the technique, while also targeting emerging post-exascale networks.
@@ 1 Introduction
Energy consumption in supercomputers and data centers is the sum of power drawn, among other elements, by computing and storage nodes, the interconnection network, and the cooling system. Considering the most powerful supercomputers, the Exascale barrier was first overrun by the Frontier supercomputer, which remained the first in the TOP500 ranking [38] since June 2022, until El Capitan surpassed it in November 2024. Table 1 contains five of the most powerful supercomputers (in terms of computing power) with energy efficiency information supplied as per the last ranking. Energy efficiency is the ratio between the maximum performance and the average power consumption; the higher it is, the better. As we can see, the highest energy efficiency is achieved by the JUPITER booster machine, which is #4 in the TOP500 ranking as of November 2025.
@@
The operational costs for supercomputers are concerning because of power availability and sustainability reasons. Green500 [37] ranks TOP500 systems according to their power efficiency rather than the raw performance. If we examine the Green500 ranking for that date, we can see that the aforementioned systems are, in terms of power efficiency, in ranks #23, #34, #90, #14 and #32, respectively. Of course, designing a system with high regard for power efficiency leads to lower operational costs, namely in electricity.
Processors and accelerators in modern supercomputers include mechanisms to reduce power under thermal or power constraints. Intel CPUs, for instance, introduced C-states [29] to enable advanced power management. These states trade performance and latency for lower power consumption and can be requested by the operating system. State transitions require microarchitectural operations and voltage changes, introducing latency, during which no code is executed. Moreover, even the deepest low-power states still consume energy. Overall, these mechanisms aim to make CPU energy consumption proportional to utilization.
@@
In this context, as computing nodes become more energy proportional, the interconnection network consumes more power, relatively [1]. The power contribution of the network can then increase, reaching up to nearly 40 % [17]. Therefore, it is also necessary to design high-performance interconnection networks whose energy consumption is somewhat proportional to the traffic they carry.
@@
Mainly, the opportunities for power savings come from periods of inactivity that would allow us to set ports to power-saving states. If we look at the network usage for several applications, we can see that network activity fluctuates during execution. Periods of low network usage are good opportunities for applying power-saving techniques. Figure 1 shows the inactivity periods for a specific port on several applications when executed on our target simulated environment. The data is then used to form a histogram of 200 fixed-sized bins. Each bin $i$ stores the number of registered inactivity periods that last between $width_{bin}\times i$ and $width_{bin}\times(i+1)$ seconds. In these plots, we show values up to the 99th percentile for visibility reasons. On the $x$ axis, we show the length of an inactivity period on that port. The left $y_{1}$ axis in black shows how many periods of inactivity are within the specified bin, and on the right, $y_{2}$ in red shows the cumulative distribution for every bin. In other words, the red line represents the percentage of the histogram’s “area” below the bin. Regardless of the distribution of inactivity periods, any power-saving mechanism should take advantage of the longer periods to provide as much energy savings as possible. Moreover, we expect the performance penalty of the mechanism to be as low as possible or within a reasonable margin.
@@
This paper explores the effects of network power management strategies on the performance of interconnection networks. First, we describe the background concepts in network energy reduction (see Section 2). Next, we describe the network power for large Ethernet-based interconnection networks used in supercomputers and data centers, and how it can be applied to BXIv3-based networks (see Section 3). We also show that more efficient interconnection networks are required to achieve energy proportionality. Afterwards, we describe the power model for BXIv3 that we have designed and the proposed energy-saving strategies. This model has been implemented in a powerful simulation framework, allowing large-scale network simulations. Furthermore, we have performed a set of simulation experiments (see Section 4) that show that, with the proposed power model strategies, the network power can be lowered without gravely affecting the overhead on execution time or latency. In particular, we managed to save more than 5 % of energy on the network by focusing on links only, while still outperforming other previous proposals in terms of performance overhead. We manage to improve the PerfBound [31] technique, even to the point of reverting energy increase into actual savings. In our opinion, this is a promising start, given that our proposal and assumptions could be portable to future interconnection network technologies, such as BXIv3 and the upcoming (Formerly Intel’s) Omnipath [^1] https://www.theregister.com/2025/06/09/omnipath_is_back/ [8].
@@
In summary, the main contributions of this paper are the following:
@@
    Study and implementation discussion of PerfBound [31], focused on the BXIv3 interconnect.
@@ 2.1 Energy consumption in high-speed interconnection networks
Network performance is affected by several network design issues, such as the topology, routing, flow control, and operating frequency, among others. Thus, designing an efficient and cost-effective network is crucial to guarantee the performance of the entire supercomputer and datacenter, as well as a balance between this performance and the investment and operational costs. Regarding HPC network technologies, a quick look into the TOP500 list shows us that Gigabit Ethernet, in various revisions, is present in many systems (34.2 %), only second to Infiniband [20] at 55.4 %. Such is the momentum of Ethernet in these environments that technologies like iWarp [10] and RoCE (RDMA over Converged Ethernet) [26] are working towards providing Ethernet with RDMA, present in interconnects like Infiniband.
@@
Beyond the substantial capital expenditure (CAPEX), operational costs (OPEX), and energy constraints are critical in HPC systems. Major OPEX components include power supply, cooling, and maintenance. Energy-saving efforts focus on reducing power consumption, potentially at the cost of performance, while cooling demands scale with heat dissipation (a direct consequence of power). At the HPC scale, both CAPEX and OPEX are enormous, and electricity costs can be decisive. Thus, reducing power consumption is a key objective, particularly given the sparse network communication in many scientific applications. In fact, InfiniBand, the most pervasive interconnect in HPC, has power-saving mechanisms [13]. The ubiquity of Ethernet in other commercial areas has also caused several proposals to appear for Ethernet as well [18, 19]. The IEEE Task Force assembled to provide the technology with a common standard, resulting in the standard IEEE 802.3az, also known as Energy Efficient Ethernet (EEE) [35].
@@
Several design and operation strategies have been applied to reduce energy consumption in interconnection networks. Specifically, network energy consumption can be reduced by designing networks with high-radix switches. By increasing the switch radix, fewer routers are required. As switches aggregate more components, this also reduces the number of hops [22]. Consequently, the number of links in the network is reduced when the hop count is lower. Routing algorithms can also be relevant to energy consumption if additional links provide several paths between two nodes. Of course, these links will consume energy when unused.
Operational costs (OPEX) in HPC systems increase with the number of nodes and network components. As with processors, network energy consumption can be reduced during periods of inactivity, which are dictated by application communication patterns. Scientific simulations typically minimize inter-node communication, resulting in sparse network activity. In contrast, data center workloads such as deep learning, distributed storage, cloud computing, and search engines exhibit network loads driven by user requests and strict latency requirements [3]. Reducing power in these multi-tenant environments is cumbersome due to certain unpredictability in network activity.
@@ 2.2 Initial attempts for reducing link power consumption
Device power consumption comprises static and dynamic components, with dynamic power depending on clock frequency, supply voltage, and transistor technology. Since power scales roughly with the cube of frequency, substantial savings are possible. Dynamic Voltage and Frequency Scaling (DVFS) adapts voltage and frequency to utilization but requires additional hardware, such as voltage regulators, increasing complexity and conflicting with the goal of simple interconnection networks [21, 33].
As network technologies often support multiple data rates, Adaptive Link Rate (ALR) was proposed to reduce energy by lowering link speeds during low utilization without changing voltage or frequency. However, rate transitions incur latency due to MAC handshakes, PHY reconfiguration, and resynchronization [18], limiting adoption and potentially creating bottlenecks in HPC workloads.
2.3 Energy-Efficient Ethernet and Low Power Idle
As ALR proved inefficient, Low Power Idle (LPI) was proposed, standardized, and used by EEE [35, 11]. LPI specifies two port power states, Wake for ports ready to transmit, and Sleep, a power-saving state, for inactive ones. This standard also specifies the basic signaling behavior, but the transition policies and implementation details are left for vendors. The LPI key point is that higher data rates are more power-efficient in terms of Joules per bit transmitted. Therefore, it would be more suitable than “downgrading” links as in ALR. The transition to Sleep consists of shutting down power-hungry PHY components so that transitions are shorter than those of ALR. However, links still draw power because they must synchronize periodically.
@@ 2.3 Energy-Efficient Ethernet and Low Power Idle
Figure 2 illustrates the interaction between packet transmissions and LPI state transitions. The parameters ${t}_{w}$ and ${t}_{s}$ denote the time required to transition from Sleep to Wake and vice versa. When packets are routed through a link in the Sleep state, the transmission can only begin after $t_{w}$, when the transition ends. Long message bursts of contiguous packets compensate for the transition overhead (one transition for the whole message), whereas short and sporadic packets may experience more frequent transition delays. These communications happen often in HPC workloads. Power consumption during state transitions is considered to be as much as in the Wake state. This damages energy saving twice, due to both the transition overhead and the high power draw.
Table 2 shows the $t_{w}$ and $t_{s}$ transition times defined by IEEE 802.3az for Ethernet rates up to 10GBase-T, along with the transmission time of a 1500-byte frame for comparison. While link speeds increase by an order of magnitude, transition times do not decrease proportionally. Still, LPI is suitable for long bursts followed by periods of inactivity comparatively longer than transmission time. Those inactive times will provide the most opportunities for power saving.
Protocol ${t}_{w}$ ${t}_{s}$ Frame time (1500B)
100Base-Tx (100 Mb/s) 30 µs 100 µs 120 µs
@@
10GBase-T (10 Gb/s) 4.48 µs 2.88 µs 1.2 µs
Table 2: Link transition times for different Ethernet versions (extracted from [35]).
Subsequent IEEE amendments refined the original LPI model to mitigate wake-up latency at higher link speeds. In particular, IEEE 802.3bj, targeting 40 Gb/s and faster Ethernet links, introduced the Fast Wake state and renamed the original low-power mode to Deep Sleep [7]. Fast Wake keeps more components of the PHY on, reducing $t_{w}$ to a tenth compared to Deep Sleep, at the cost of reduced power savings (20–40 % compared to up to 90 % in Deep Sleep[30]. Figure 3 conceptually illustrates the difference between these low-power states.
@@
These multiple power levels provide the basis for energy-aware link management strategies in Ethernet networks. In HPC environments, several proposals leverage Fast Wake to limit the latency impact of LPI [28, 2]. However, transitioning to low-power states instantly after a packet transmission may cause links to spend more time in state transitions than in active data transmission, reducing the potential energy savings.
IEEE specifications beyond 40GBase-T do not explicitly report LPI transition times, which depend on the PHY implementation and medium. For higher-speed links, we estimate wake and sleep transition durations using a power-law extrapolation based on measured PHY generations. Denoting the link rate as $R$, the wake-up time is modeled as $t_{w}\propto R^{b}$, yielding $\log(t_{w})=b\,\log(R)+c$. A least-squares fit in log–log space results in a scaling exponent of b=0.45b=-0.45 for wake transitions and b=0.77b=-0.77 for sleep transitions.
Table 3 summarizes the measured LPI transition times up to 10GBase-T and the extrapolated values for higher Ethernet rates. These estimates represent indicative trends rather than standard-compliant parameters.
Ethernet rate $\boldsymbol{t_{w}}$ (µs) $\boldsymbol{t_{s}}$ (µs)
100Base-TX 30 (measured) 100 (measured)
1000Base-T 16 (measured) 182 (measured)
10GBase-T 4.48 (measured) 2.88 (measured)
40G Ethernet 2.4 (estimated) 1.2 (estimated)
100G Ethernet 1.6 (estimated) 0.7 (estimated)
400G Ethernet 0.8 (estimated) 0.25 (estimated)
Table 3: Measured and extrapolated LPI wake ($t_{w}$) and sleep ($t_{s}$) transition times as a function of Ethernet bandwidth.